Yui101:

Dillon can buy a fixed-rate bond with a 4% coupon, a $10,000 principal, and five years left to maturity or a TIPS with the same principal and age to maturity with a 2% coupon. Compare the two bonds if inflation runs 3% per year on average. The fixed-rate bond has a better yield. The TIPS has a better yield. Both yields are roughly equivalent. The yields vary with inflation.

AnimeGhoul8863:

WELCOME TO QUESTIONCOVE! YUI!!

Yui101:

thank you ^^ happy to be here

AnimeGhoul8863:

give me a few minutes and ill have your answer

Yui101:

ok thank you no rush ^^

Yui101:

Btw AnimeGhoul8863 i love your profile pic she is from the dragon maid i forgot what its called but she is so cute ^^

AnimeGhoul8863:

LOL its kanna from miss kobayashis dragon maids and you can call me Anime or Ghoul which ever you prefer

Yui101:

lol ok thank you and yeah her its been awhile since ive seen it ive completed the serious really good ^^ and may i call you Anime

AnimeGhoul8863:

Sure lol and if you have any picture and wnat your name added to it with hearts or stuff talk to meh and i can make it

Yui101:

alright ^^ i will thank you

AnimeGhoul8863:

@Elsa213

AnimeGhoul8863:

srry i got nothing maybe Elsa can help

Yui101:

ok ^^

Elsa213:

@Shadow e.e

Elsa213:

Erm... Whale, I googled. e.e I found something. o; ``` tkhunny 4 years ago Sounds like a job for a present value. 10000*0.04/2 = 10000*0.02 = 200 -- Okay, there's the one coupon. Go! ```

Yui101:

thank you Elsa for helping ^^ im srry if its a hard one

Elsa213:

Math isn't my best subject, sorry.

Yui101:

its ok its not mine either im almost done with it i just how two things to do and ill get my full credit for math models and no more math for me im almost close to graduating

Elsa213:

Congratulations. :D

Yui101:

thank you ^^

Yui101:

imma close this post bc i just guessed lol bc i have a new question